OpenStudy (anonymous):

Two trains leave the station at the same time, one heading west and the other east. The westbound train travels at 90 miles per hour. The eastbound train travels at 100 miles per hour. How long will it take for the two trains to be 456 miles apart?

OpenStudy (australopithecus):

 well they both start at 0 miles at time = 0, so you can probably just add their speeds to figure out the speed at which they are moving apart 90miles per hour + 100miles per hour = 190miles per hour 190mile/h is the speed at which they move apart set up equality and solve for x 190mile/1 hour = 456mile/x hour

OpenStudy (anonymous):

I didn't get it >.<

OpenStudy (australopithecus):

do you understand how I got 190 miles per hour?

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

ok 1 sec

OpenStudy (poopsiedoodle):

I dunno, but whoever is observing the trains must have no life.

jimthompson5910 (jim_thompson5910):

Another way to do it: westbound train: distance = rate * time d = rt d1 = 90t eastbound train: distance = rate * time d = rt d2 = 100t ---------------------------- To find out how long it will take them to be 456 miles apart, add the two distances and set it equal to 456 d1 + d2 = 456 90t + 100t = 456 190t = 456 t = 456/190 t = 2.4 So it will take 2.4 hours (2 hours, 24 minutes)

OpenStudy (anonymous):

@zello Check your previous post! It takes the solver hours to explain it to you, then you don't know ANY at all !!!

jimthompson5910 (jim_thompson5910):

doing it Australopithecus's way will give you the same answer

OpenStudy (anonymous):

It's different

OpenStudy (anonymous):

Yeah d1 + d2! That's it :)

OpenStudy (anonymous):

So how do i now if it's d1+d2 or d1=d2

OpenStudy (australopithecus):

so you set up a ratio so the units cancel (kind of like numbers